OpenStudy (anonymous):

Write a rule for the sequence. Then find the next three terms. 6, 8, 10, 12, . . .

OpenStudy (hunus):

The numbers are increasing by a fixed amount for each new number.

OpenStudy (ja1):

well they are going up by 2 (+2)

OpenStudy (ja1):

so if you add 12 +2 =

OpenStudy (anonymous):

=14

OpenStudy (ja1):

yes and so on 14 +2 = 16

OpenStudy (ja1):

16 + 2 = ?

OpenStudy (dan815):